OpenStudy (anonymous):

what is the solution of the matrix equation? [ 5 -2] X= [2] 2 -1 -4

OpenStudy (anonymous):

So you have any idea how to find the order of matrix X ?

OpenStudy (anonymous):

no clue..

OpenStudy (anonymous):

Ok at least you know that order of [ 5 -2 2 -1 ] is (2x2)?

OpenStudy (anonymous):

yea i do

OpenStudy (anonymous):

Yes. so let Order of matrix X be (a x b ) And order of product matrix is (2x1) We know that matrices can be multiplied if no of columns of 1st matrix equals no. of rows of X. Then no . of rows of X = 2. And if (a x b) . ( c x d) matrix is multiplies, result is always matrix of order (a x d) Hence no. of columns in X = 1. (since product matrix of order (2x1) would be d= 1) So order of X = ( 2 x 1 ). Ok till here?

OpenStudy (anonymous):

ya i got it so far i think

OpenStudy (anonymous):

[ 5 -2] ( x y ) = [2] 2 -1 -4 This would mean - 5.x - 2.y => 2 2.x - 1.y -4 (5x - 2y => 2 2x - y) -4. Now the matrices are equal; hence their corresponding elements are equal as well. so 5x-2y = 2, 2x-y = -4, ok?

OpenStudy (anonymous):

alright ya. so now how do you solve it from here?. it's not like a linear equation is it?

OpenStudy (anonymous):

Well nope. these are simultaneous equations. 5x-2y = 2, 2x-y = -4. You can solve it using substitution method. 2x - y = -4 so y = 2x+4 Use y = 2x + 4 in 5x-2y = 2, u'll get the value of x. hence solve the value of Y. The matrix X will be (x y)

OpenStudy (anonymous):

oh alright i didnt think it was linear, im just not familiar with matrices

OpenStudy (anonymous):

haha no probs..hope u got the right ans :)

OpenStudy (anonymous):

me too thanks for the help!